### v1

# Token Dynamics: Towards Efficient and Dynamic Video Token Representation for Video Large Language Models

Abstract
Token-based video representation has emerged as a promising approach for enabling large language models (LLMs) to interpret video content. However, existing token reduction techniques, such as token pruning and token merging, often disrupt essential spatial-temporal positional embeddings, failing to adequately balance computational efficiency with fewer tokens. Consequently, these methods result in relatively lengthy token sequences, limiting their applicability in scenarios requiring extreme token compression, such as video large language models. In this paper, we introduce the novel task of extreme short token reduction, aiming to represent extensive video sequences with a minimal number of tokens. To address this challenge, we propose Token Dynamics, a new video representation framework that dynamically reduces token count while preserving spatial-temporal coherence. Specifically, we disentangle video representations by separating visual embeddings from grid-level motion information, structuring them into: (i) a concise token base, created by clustering tokens that describe object-level content, and (ii) a token dynamics map, capturing detailed spatial-temporal motion patterns across grids. Furthermore, we introduce a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, thereby maintaining compactness and spatial-temporal integrity. We evaluate Token Dynamics on the proposed extreme short token reduction task, demonstrating a reduction of token count to merely 0.07% of the original tokens, with only a minor performance drop of 1.13% on the NextQA-MC benchmark. Our approach similarly maintains competitive performance on ActNet-QA, Long Video Bench, and VideoMME datasets in a plug-and-play manner. Additionally, we propose two novel subtasks within extreme token reduction (fixed-length and adaptive-length compression) both effectively representing long token sequences for video-language tasks. Our method achieves state-of-the-art performance across these subtasks, offering significantly lower theoretical complexity, fewer tokens, and enhanced throughput, thus providing an efficient solution for video large language models.
1 Introduction
[…]
Baseline (LLaVA-OV) 11664 100% 46 21.91 8.2s 0 $\mathcal{O}(n^{2}dL)$ 58.38
Token Pruning ($k=0.9n$) 1152 10% 89 16.09 4.3s $\mathcal{O}(n^{2}d)$ $\mathcal{O}((n-k)^{2}dL)$ 29.12
ToMe ($k=0.9n$) 1152 10% 42 11.53 9.0s $\mathcal{O}(n^{2}d)$ $\mathcal{O}((n-k)^{2}dL)$ 35.72
VidToMe ($k=0.9n$) 1152 10% 40 11.49 9.4s $\mathcal{O}(n^{2}d)$ $\mathcal{O}((n-k)^{2}dL)$ 39.64
Interpolating ($k=0.73n$) 3136 27% 32 13.59 11.8s $\mathcal{O}(nd)$ $\mathcal{O}((n-k)^{2}dL)$ 57.20
Ours-Dynamic ($m$ adaptive) 16 0.07% 49 10.50 7.8s $\mathcal{O}((n+m^{2})d)$ $\mathcal{O}(m^{2}dL)$ 57.72
Ours-Fixed ($m=32$) 32 0.14% 91 10.47 4.2s $\mathcal{O}((n+m^{2})d)$ $\mathcal{O}(m^{2}dL)$ 57.46
Table 1: Comparison of model efficiency in terms of token number, throughput, FLOPs, run-time, accuracy, and complexity analysis. Note: $n$ is the original token count; $k$ is the number of tokens reduced by traditional methods; $m$ is the compressed token count after our extreme reduction approach, with the relationship $n>k\gg m^{2}\gg m$; $d$ denotes token dimensionality; and $L$ represents transformer layer count. Given $m\ll n$, our token reduction module has a complexity of $\mathcal{O}((n+m^{2})d)\approx\mathcal{O}(nd)$, significantly reducing LLM complexity to $\mathcal{O}(m^{2}dL)$. Module Complexity quantifies the computational cost of the token reduction method itself, while LLM Complexity reflects the computational reduction within the LLM, benefiting from the reduced token count.
Recent advances in Vision Language Models (VLMs) have enabled unified zero-shot capabilities across diverse tasks, including visual question answering [25, 12], video-to-text generation [1], video segmentation [26], and video understanding [29]. Although VLMs excel at aligning visual and linguistic information, their substantial computational cost remains a critical bottleneck—especially for video large language models (vLLMs). Video inputs contain spatial-temporal information distributed across numerous frames, resulting in lengthy token sequences that significantly burden computational resources [6, 27]. Consequently, as vLLMs scale in size [12, 29], improving computational efficiency becomes imperative.
Unlike textual inputs, video data require tokenizing pixel batches from each frame and concatenating them into extensive sequences. Transformers process these sequences through attention mechanisms at each layer, incurring a computational complexity of $\mathcal{O}(n^{2}DL)$. As demonstrated in Table 1, the token sequence length ($n$) is the primary contributor to computational overhead, increasing exponentially as the token count grows. This overhead surpasses the influence of model parameters, layers ($L$), and embedding dimensions ($D$). Reducing token sequence length emerges as a promising solution, broadly applicable to most LLMs in a plug-and-play manner.
Despite extensive research on token reduction techniques to remove redundant information across video frames, existing approaches encounter three primary challenges. Token pruning methods [10, 17] aim to eliminate redundant tokens but frequently discard critical information, impairing accurate representation. Token merging strategies such as ToMe [3], Vid-ToMe [11], and Token Bilinear Interpolating [12] merge similar tokens to improve efficiency without explicitly discarding them. However, these methods typically employ fixed token reduction ratios, limiting flexibility and leaving sequences excessively long, particularly for large-scale video data.
We attribute these challenges to the lack of dynamic strategies for token selection and compression. Existing methods relying on fixed-number or fixed-percentage strategies yield suboptimal outcomes: excessively long sequences retain redundant information, while overly short sequences risk severe performance degradation. Furthermore, current approaches lack adaptive context sensitivity for optimal token selection. To address these limitations, we propose a dynamic token reduction strategy based on adaptive clustering using grid-level similarity, enabling dynamic grouping of similar tokens into concise sequences. However, preserving motion dynamics during clustering remains challenging.
The second major challenge involves maintaining spatial-temporal relationships during token reduction. Traditional token pruning methods [10, 17] frequently disrupt positional information crucial for accurately tracking object movements and preserving motion continuity. Similarity-based merging techniques [3, 11, 12] often disregard spatial-temporal coherence by either directly merging positional encodings or reintroducing them post-merging, which impairs effective dynamic context representation.
To address this, we introduce a structured token map representation, organized as Temporal $\times$ Width Coordinates $\times$ Height Dimension, to explicitly record spatial-temporal positions before clustering. To incorporate these dense maps without additional computational overhead, we propose a cross-dynamics attention mechanism that embeds motion information into clustered token sequences while maintaining their compact structure. Given video data’s inherent redundancy, tokens effectively represent grid-level visual information derived from pixels, offering a lightweight alternative to computationally intensive motion-tracking methods such as optical flow. Tracking token movements through the token map efficiently captures object grid-level dynamics, preserving essential spatial-temporal context with minimal computational cost.
The third challenge lies in effectively evaluating concise token sequences. The limitations of existing methods prevent significant token reduction, making direct comparisons of token-performance trade-offs and adaptability difficult. We define the extreme token reduction task, comprising two subtasks: (i) fixed-length token compression for evaluating token-performance efficiency, and (ii) adaptive-length compression to assess adaptability. Additionally, we introduce separate complexity metrics for the token reduction module itself and its impact on LLM complexity, providing a comprehensive evaluation framework for extreme token reduction methods.
Our contributions are summarized as follows:
  1. (i)
    We introduce a novel video representation framework featuring fixed/adaptive token clustering. It constructs a compact token base to capture visual pattern and object part descriptions and a token dynamics map to model their movements. A cross-motion dynamics attention mechanism integrates spatial-temporal positional information without increasing token count.
  2. (ii)
    We define the extreme token reduction task, aiming to represent videos with minimal tokens. This includes fixed-length and adaptive-length subtasks, along with new efficiency metrics that effectively distinguish between module complexity and LLM processing demands.
  3. (iii)
    Our method reduces token count to just 0.07% of the original sequence with minimal performance loss. It significantly outperforms existing approaches in token reduction efficiency under both adaptive and fixed-number settings, demonstrating effectiveness and better performance trade-offs.
2 Related Works
2.1 Video Large Language Models
Video Large Language Models (vLLMs) have emerged as powerful tools for bridging video understanding and natural language processing, enabling complex interpretations of video content through language-based interactions. Recent advancements have demonstrated remarkable capabilities in aligning visual and linguistic modalities, exemplified by frameworks such as LLaVA [15, 12, 14, 16], Flamingo [1], and MERLOT Reserve [29]. These methods typically rely on extensive pre-training using large-scale datasets like HD-VILA [26], InternVid [23], and NextQA [25], generating lengthy token sequences to represent videos effectively.
2.2 Token Reduction
Token reduction techniques have gained increasing attention as a means to enhance computational efficiency in Vision Transformers (ViTs) [6]. Early influential methods, such as Token Pruning [10] and Token Merging (ToMe) [3], significantly reduce computational load by identifying and removing redundant tokens or merging similar tokens, respectively. More recently, Video Token Merging (Vid-ToMe) [11] has extended ToMe to video inputs, considering temporal redundancy by merging tokens across consecutive frames. Despite these advancements, existing token reduction strategies generally adopt fixed token reduction ratios (e.g., 50%), limiting flexibility and adaptability. Such fixed strategies can either inadequately reduce redundant tokens, resulting in lingering inefficiencies, or inadvertently merge tokens representing distinct objects, thus losing crucial spatial-temporal dynamics necessary for precise video interpretation. To overcome these limitations, our proposed Token Dynamics framework introduces dynamic token clustering to generate a compact token representation while explicitly preserving spatial-temporal motion information via a dedicated token dynamics map. Through our novel cross-dynamics attention mechanism, our approach effectively integrates spatial-temporal coherence into concise token sequences without compromising accuracy, outperforming state-of-the-art token reduction methods, even in scenarios of extreme token compression.
Figure 1: Overview of the Token Dynamics framework and its integration into video large language models. The input video is first tokenized into a 1D token sequence and subsequently undergoes token clustering to group similar tokens, potentially representing the same object parts at the grid level. This clustering yields our proposed Token Dynamics representation, comprising two complementary components: (i) a concise token base, dynamically sized by merging tokens into representative cluster centroids, and (ii) a token dynamics map, a three-dimensional tensor explicitly capturing grid-level spatial-temporal coordinates (i.e., x, y, and frame index as spatial-temporal positions across frames) to represent detailed motion dynamics for each cluster. To efficiently integrate motion information into the token base, the token dynamics map is first flattened and dimensionally aligned via an MLP, and then processed by a cross-dynamics attention module employing two multi-head attention operations. Crucially, this integration preserves the compactness of the original token count. Finally, the compressed token sequence, along with a tokenized query, is projected into the large language model for zero-shot inference.
3 Methods
3.1 Problem Definition: Extreme Token Reduction
The primary goal of the Extreme Token Reduction task is to develop a method that represents a long token sequence $t$, typically provided to a large language model (LLM), using only a minimal number of tokens, while preserving comparable performance.
Formally, given a video $v$ and a query question $q$ as inputs, where $q$ serves as a prompt to extract specific information from the video, the task for a classical video large language model (vLLM) is to predict the corresponding answer $a$. To facilitate this, the video $v$ is first tokenized into a sequence $t$. The token reduction method takes this original long token sequence $t$, derived from video frames, and outputs a significantly shorter yet sufficiently informative token sequence. The extreme token reduction task consists of two subtasks, alongside a new metric for separate analyses of computational complexity introduced by the token reduction module and its impact on subsequent LLM.
3.1.1 Fixed-Length and Adaptive-Length Token Reduction Subtasks
The fixed-length token reduction subtask compares the performance of different token reduction methods under a predefined token reduction rate. This allows fair benchmarking since differing token lengths can significantly influence performance, complicating comparisons. Current methods usually require an explicit token merging or pruning rate, making this subtask particularly suitable for evaluating these approaches.
However, the flexibility to dynamically adjust token count according to input complexity is critical for practical applications. Different input videos inherently require varying token numbers to adequately capture essential information. Therefore, we define an additional subtask, adaptive-length token reduction, where methods dynamically determine an optimal token count tailored to each input video, enabling a fairer assessment of their adaptability and effectiveness.
3.1.2 Metrics: Separating Module Complexity and Subsequent LLM Complexity
Computational complexity is conventionally evaluated by aggregating the complexity of all components within a pipeline. However, token reduction techniques offer dual impacts: they introduce additional complexity themselves but subsequently alleviate computational burdens by significantly reducing the token count processed by the LLM.
To better characterize these distinct impacts, we introduce two separate metrics:
  • Module Complexity: Evaluates the computational cost of all operations introduced specifically by the token reduction module itself.
  • LLM Complexity: Quantifies the computational efficiency gains achieved by the reduced token count when processed by the LLM, typically represented by the significantly lowered complexity due to shorter token sequences.
These metrics clearly disentangle the direct computational overhead introduced by token reduction modules from the substantial computational benefits realized through reduced token processing in LLMs.
3.2 Preliminaries: Token Merging (ToMe)
Token Merging (ToMe) [3] improves the efficiency of Vision Transformers (ViTs) [6] by merging similar tokens within transformer layers, reducing computational cost without requiring additional training. The merging operation is formally defined as:
$t_{i}=\mathcal{M}(t,j)=\frac{\sum_{j}m_{j}t_{j}}{\sum_{j}m_{j}},$
where $\mathcal{M}$ represents the merging function, $t_{j}$ is the $j$-th token within a group of similar tokens, $m_{j}$ is the corresponding weighting factor, and $t_{i}$ is the merged token.
ToMe applies a fixed merging ratio, which can lead to suboptimal token aggregation. A small ratio results in incomplete merging, leaving redundant tokens, while a large ratio risks merging tokens from distinct objects, disrupting object coherence. Additionally, ToMe alters spatial-temporal positional encodings, which may degrade video representation quality. These limitations highlight the need for an adaptive approach that better preserves object-level semantics and spatial-temporal consistency.
3.3 Token Dynamics Representation
We introduce the Token Dynamics Representation, designed to minimize token sequence length dynamically while preserving complete spatial-temporal motion information. We conceptualize videos as compositions of multiple scenes, each containing objects with distinct attributes and motions. To achieve this, we reformulate the original video representation (token sequences from Vision Transformers [6]) into two components: (i) a concise token base, representing objects in each scene as token centroids obtained by clustering similar visual patterns at the grid-level; (ii) a token dynamics map, explicitly recording the spatial-temporal coordinates of each cluster centroid, capturing detailed object-level grid trajectories. As illustrated in Fig. 1, this design ensures the token base and dynamics map are disentangled at the grid-level. Thus, we represent video content concisely without sacrificing crucial spatial-temporal information, achieving efficiency in token sequence reduction.
3.3.1 Token Clustering
Token sequences derived from Vision Transformers (ViTs) [6] often exhibit redundancy, as visual patterns evolve over time but frequently represent the same objects. While tokens capture semantic information, slight variations necessitate clustering to effectively group similar patterns across frames.
Previous methods, such as ToMe [3], rely on fixed merging ratios, risking incomplete merging or incorrect grouping of tokens. To mitigate these issues, we adopt classical K-Means [22] for fixed-length clustering and introduce an adaptive K-Means [2] variant for dynamic token selection. Token similarity is computed using cosine similarity:
$(s_{0},s_{1},\dots,s_{K}),(c_{1},c_{2},\dots,c_{N})=\mathcal{F}_{\text{cluster}}(t_{1},t_{2},\dots,t_{N}),\quad C_{i}\in[1,K]$
where $K$ is the number of clusters (fixed or adaptive), $\mathcal{F}_{\text{cluster}}$ denotes the clustering function (e.g., K-Means, Adaptive K-Means, or Video Segmentation [20]), $t_{i}$ represents the $i$-th token, $c_{i}$ is its assigned cluster index, and $s_{k}$ is the set of token indices assigned to the $k$-th cluster.
3.3.2 Concise Token Base
To construct a compact representation while preserving essential visual patterns and object part information, we merge clustered tokens into representative centroids, forming the concise token base. Using the token indices from Equation 2, the $k$-th token base $b_{k}$ is computed as:
$b_{k}=\mathcal{M}(t,s_{k}),\quad k\in[1,K],\quad b\in\mathbb{R}^{K\times D}$
where $b_{k}$ represents the centroid token for the $k$-th cluster, $s_{k}$ is the set of token indices assigned to the $k$-th cluster, and $\mathcal{M}$ denotes the merging function from Equation 1.
3.3.3 Token Dynamics Map
While the concise token base effectively represents visual patterns, it lacks spatial-temporal details essential for video understanding. To preserve motion dynamics, we record spatial-temporal coordinates during clustering, constructing a detailed motion representation—i.e., the token dynamics map.
Specifically, we formulate a three-dimensional tensor capturing $x$, $y$, and frame indices, ensuring that each cluster centroid retains its original spatial-temporal position. This map aligns with the downsampled video structure from the ViT and is defined as:
$m_{f,x,y}=c_{(fWH+xW+y)},\quad m\in\mathbb{R}^{T\times W\times H}$ (4)
where $T$, $W$, and $H$ denote the frame count, width, and height of the tokenized video, respectively; $c_{i}$ is the cluster index of the $i$-th token; and $m_{f,x,y}$ represents the token map entry at frame $f$ and spatial coordinates $(x,y)$.
3.4 Cross Dynamics Attention Module
While the token base and token dynamics map effectively disentangle visual and motion information, directly embedding the token dynamics map would significantly increase computational overhead due to its large size. To address this, we propose a cross-dynamics attention module that integrates motion dynamics into the concise token base without increasing token count or dimensionality.
First, the token dynamics map is projected via an MLP to align its dimensions with the token base. Then, a multi-head transformer decoder extracts motion information by treating the projected token dynamics map as encoded features and the token base as queries. This process embeds spatial-temporal dynamics into the token sequence while maintaining its compact structure:
$b_{\text{bank}}^{K\times D}=\mathcal{F}_{\text{A}}^{K\times D}\big{(}m^{T\times W\times H}W_{1}^{H\times D},\mathcal{F}_{\text{A}}^{K\times D}(b^{K\times D}W_{2}^{D\times D})\big{)}$
where $\mathcal{F}_{\text{A}}$ represents multi-head attention, $m$ is the token dynamics map, $b$ is the concise token base, and $W_{r}$ denotes the weight matrix in the $r$-th MLP layer. Superscripts indicate tensor shapes. The output $b_{\text{bank}}$ is the refined compressed token sequence, which is subsequently fed into the LLM for downstream tasks.
3.5 Plug-and-Play Integration with vLLMs
Our Token Dynamics framework, combined with the cross-dynamics attention module, is designed as a plug-and-play solution for video large language models (vLLMs). It compresses token sequences from the visual tokenizer, integrates motion dynamics, and feeds the processed sequences directly into the original projection layer of the vLLM. The query question tokens are concatenated with the projected compressed token sequence, and the entire framework is trained end-to-end using the original vLLM’s loss function.
4 Experiments
4.1 Implementation Details
4.1.1 Datasets
Training Dataset. We follow the training setup of LLaVA-OneVision [12], containing 178K videos paired with 1.3M instruction-following samples. These include 178K captions, 960K open-ended questions, and 196K multiple-choice questions, ensuring balanced coverage of diverse video scenarios.
Evaluation Benchmarks. To comprehensively evaluate our model’s effectiveness across diverse scenarios, we select six benchmarks that vary in video length, resolution, and task complexity, covering spatio-temporal reasoning, long-form understanding, and multimodal comprehension. ActivityNet-QA [28] (videos up to 120s) assesses spatio-temporal reasoning in web videos, while MLVU [31] (3 min – 2h) focuses on multimodal video understanding across movies, TV shows, and documentaries. VideoMME [7] average 17 min provides a real-world long-video understanding challenge. NExT-QA [25] examines descriptive, causal, and temporal reasoning. LongVideoBench [24] (up to 1h) targets long-term narrative and event understanding, and MVBench [13] (35s videos) introduces 20 reasoning-intensive tasks requiring temporal and conceptual comprehension. These benchmarks ensure a comprehensive evaluation of our approach across varied durations, resolutions, and complexities to validate robustness.
4.1.2 Metrics
Accuracy (Acc.): Accuracy measures the percentage of correct responses generated by the model in multiple-choice and open-ended video question-answering tasks, reflecting its effectiveness in understanding video content. Token Count (Token Num.): The number of tokens processed by the model per task, serving as an indicator of computational efficiency. Lower token counts imply better compression and more concise representations while preserving information. Throughput: Measured as images processed per second, calculated as the total images processed divided time in seconds. FLOPs (T): Total floating-point operations required for inference, indicating computational cost. Run-Time: Overall inference time, capturing the efficiency of the model in processing video content. Module Complexity: Time complexity overhead introduced by the token reduction module itself, quantifying its additional computational cost. LLM Complexity: The computational complexity of the large language model after incorporating the token reduction module, primarily influenced by the reduced token sequence length.
4.1.3 Implementation Details
Our model is trained on the LLaVA-OneVision dataset and evaluated in a zero-shot setting on multiple benchmarks. To ensure practical efficiency, we adopt the 0.5B-parameter LLaVA-OneVision [12], based on QWen2 [27], as our baseline. We integrate our proposed Token Dynamics framework into this baseline and fine-tune it for evaluation. Training is conducted on 4 NVIDIA A100 GPUs for 85,000 iterations, using the AdamW optimizer. The learning rate follows a cosine decay schedule, starting at $1\times 10^{-5}$ for the cross-dynamics attention module and $2\times 10^{-6}$ for the ViT module. We employ Zero2 [19] optimization with a batch size of 8 and gradient accumulation steps of 2.
4.1.4 Video Large Language Model Baselines
Due to limitation of computation resources, we compare the vLLMs on 0.5B parameters track. As of now, the only available 0.5B video large language model baselines are LLaVA-OV-SI and LLaVA-OneVision. For efficiency, we plug our model into the LLaVA-OneVision baseline for quick training and evaluation, minimizing GPU consumption. To ensure a fair comparison, we evaluate models separately for the 0.5B and 7B parameter configurations. While 7B models have 14 times the parameters and benefit from significantly more GPU resources, our 0.5B model—with only 0.14% of token numbers—outperforms some 7B models, demonstrating our method’s efficiency and effectiveness.
4.1.5 Token Reduction Baselines
For token reduction, we compare against several baselines: Token Pruning [10]: A widely recognized method for reducing token numbers and increasing throughput in LLMs.
Token Merging (ToMe) [3]: A popular baseline for token reduction, known for its efficiency improvements.
Video Token Merging [11]: The current state-of-the-art method for token reduction in video large language models, extending the capabilities of ToMe to video data.
Interpolation: Introduced by [12], the use of bilinear interpolation to reduce the number of tokens in visual representations, particularly for video frames. This approach allows the model to handle a larger number of frames by reducing the tokens per frame, achieving a balance between performance and computational cost.
Table 2: Zero-shot performance comparison of video large language models. We compare the performance of our Token Dynamics model with recent video large language models (2022–2024) on video understanding tasks. Starting from the 0.5B LLaVA-OneVision baseline, we integrate Token Dynamics to achieve significant token reduction. Our model demonstrates competitive accuracy with only a 1.6% drop while using just 0.14% of the original token numbers. Furthermore, it outperforms other 0.5B models and even some 7B models.
Model #Parameters Year Zero-Shot Acc.(%) $\uparrow$ Token Num.$\downarrow$
Mistral [8] 7B 2023 51.1 100%
P3D-G [5] 7B 2022 51.33 100%
VFC [18] 7B 2023 51.5 100%
LLoVi [30] 7B 2023 54.3 100%
MVU [21] 7B 2024 55.2 100%
ATP [4] 7B 2022 54.3 100%
LLaVA-OneVision [12] 0.5B 2024 57.2 100%
LLaVA-OV-SI [12] 0.5B 2024 53.6 27%
Token Dynamics (Ours) 0.5B 2024 57.46 0.14%
4.2 Quantitative Comparison with vLLM Baselines
Table 2 presents a comparison of our Token Dynamics model against recent video large language model (vLLM) baselines introduced between 2022 and 2024. To ensure a fair evaluation, we categorize baselines based on model size—0.5B and 7B parameters—accounting for the substantial performance differences driven by model scaling, as described by the neural scaling law [9]. Although our model was trained separately and evaluated in a strict zero-shot setting, we also include non-zero-shot baselines trained directly on evaluation datasets for a more comprehensive assessment. We evaluate all models based on two primary metrics: Accuracy and Token Count. Our Token Dynamics model slightly outperforms the LLaVA-OneVision baseline in accuracy while drastically reducing token numbers. The LLaVA-OneVision model processes 23,328 tokens (100%), whereas our model achieves a comparable accuracy while utilizing only 32 tokens (0.14%), significantly lowering computational costs. Moreover, despite its smaller size (0.5B parameters), our model surpasses several recent 7B vLLMs in zero-shot accuracy, demonstrating its effectiveness in token reduction without sacrificing performance. These results highlight the strength of our Token Dynamics framework in efficiently eliminating redundancy while preserving essential spatial-temporal and semantic information. The drastic token compression achieved by our model underscores its potential for enhancing the efficiency of video-language models, making large-scale video understanding more computationally feasible.
4.3 Token Reduction Task
4.3.1 Fixed-Length Subtask
While our previous experiments highlight the effectiveness of our model in reducing token numbers while maintaining accuracy compared to recent vLLMs, it is also crucial to benchmark its efficiency against classical and state-of-the-art token reduction methods, such as Token Pruning, Token Merging (ToMe), and Video Token Merging (VidToMe). These approaches typically apply a fixed token reduction ratio (e.g., 50%). However, since token reduction significantly influences LLM throughput and efficiency, evaluating these methods under extreme compression conditions provides deeper insight into their effectiveness. To ensure a fair comparison, we configure our Token Dynamics model with a fixed clustering number $K$, allowing us to control the exact number of tokens retained. We then set each baseline method to reduce tokens to the same predefined lengths (e.g., 12, 32, or 64 tokens), ensuring all models operate with identical input token counts. This standardized evaluation isolates the impact of each reduction strategy on model accuracy. As shown in Table 3, our model consistently outperforms Token Merging (applied per frame), Video Token Merging (applied across entire video sequences), and Token Pruning (which discards the least important tokens). This superior performance highlights the effectiveness of our Token Dynamics framework in achieving extreme token reduction while preserving accuracy, setting a new benchmark for efficiency in video-language models.
4.3.2 Adaptive Length Subtask
In this subtask, each method must be capable of adaptively determining the optimal number of tokens based on the input token sequences. Unlike fixed-length approaches, adaptive methods dynamically adjust token counts to balance efficiency and performance. As shown in Table 4, we compare the performance of different baselines and our model in adaptive token length selection. The ”Avg Tokens” column represents the average number of tokens used across all test samples for each method. Compared to the token interpolation baseline [12] and our fixed-length approach, our adaptive model achieves higher accuracy while consuming significantly fewer tokens on average.
Table 3: Comparison of Fixed-Length Token Reduction Methods. We evaluate different token reduction approaches by setting a fixed number of retained tokens. Each baseline method is adjusted to match the same token count for a fair comparison of accuracy and computational efficiency.
Baseline Preset Length Acc.$\uparrow$ 12 32 64
Token Pruning 29.12 34.50 31.31
ToMe 35.72 38.50 40.10
VidToMe 39.64 45.10 46.20
Ours (Fixed) 57.03 57.46 57.10
Table 4: Comparison on Adaptive Length Token Reduction. Models select token lengths dynamically based on input sequences. For baselines, we use their default settings optimized for most cases to ensure a fair comparison.
[…]
Interpolating [12] 3136 57.20
Ours (Fixed) 32 57.46
Ours (Dynamic) 13.08 57.72
4.4 Efficiency Comparison and Analysis
To comprehensively evaluate the practical efficiency gains from different token reduction methods, we compare our approach against existing techniques using appropriately chosen token reduction settings. For Token Pruning, ToMe, and VidToMe, we retain 10% of the original tokens, a commonly used configuration balancing efficiency and performance. For Token Interpolating, we use its default bilinear interpolation setting, which retains 27% of the original tokens. For our Token Dynamics approach, we configure Ours-Fixed to use the best-performing token count from Table 3, while Ours-Dynamic adaptively determines token numbers using an adaptive K-Means algorithm [2]. We evaluate efficiency across seven key metrics: Token Number, Token Number Percentage, Throughput (measured in images per second), FLOPs, Run-Time, Module Complexity (overhead introduced by the token reduction module), and LLM Complexity (computational reduction achieved within the LLM due to token number reduction). To ensure a fair comparison, we use vanilla LLaVA-OV as the baseline and apply each token reduction method to analyze its impact on efficiency and performance.
As shown in Table 1, our Ours-Fixed and Ours-Dynamic approaches outperform all baselines across efficiency and accuracy metrics, particularly in theoretical time complexity analysis. These results highlight the effectiveness of Token Dynamics in achieving a superior trade-off between token reduction and accuracy, demonstrating state-of-the-art efficiency without compromising performance.
4.5 Performance on Multiple Benchmarks
Evaluation Across Diverse Resolutions and Complex Scenarios. To assess the effectiveness of token reduction methods in real-world video settings—including high-resolution, long-form datasets and complex tasks such as multi-step reasoning—we extend our evaluation to additional benchmarks (introduced in Sec. 4.1.1). These benchmarks encompass a variety of video types, resolutions, and reasoning complexities. As shown in Table 5, our Token Dynamics approach achieves comparable accuracy to baselines while using 99.86% fewer tokens. This highlights the robustness of our method in maintaining performance even under extreme token reduction.
Benchmark Token Num.(%) NextQA-MC ActNet-QA Long Video Bench VideoMME
LLaVA-OV-SI 100% 53.6 49.0 41.9 40.4
LLaVA-OneVision 100% 57.2 50.5 45.8 43.5
Token Dynamics (Ours) 0.14% 57.44 46.25 39.34 38.15
Table 5: Performance comparison across multiple benchmarks. Our method significantly reduces token numbers while maintaining competitive accuracy across diverse video understanding tasks.
Evaluation Across Multiple Subtasks. To further evaluate the effectiveness of Token Dynamics, we assessed its performance on 20 diverse subtasks from MVBench, spanning pose estimation, navigation, multi-step reasoning, and object interactions in highly dynamic video environments. As shown in Fig. 2, our model consistently achieves competitive performance across most subtasks. Notably, it demonstrates superior accuracy in action and object interaction tasks, highlighting its ability to enhance the model’s focus on dynamic object relationships and motion understanding.
Figure 2: Performance comparison across multiple MVBench subtasks. Our method consistently performs well across diverse reasoning and interaction-based tasks, particularly excelling in action recognition and object interaction.
VLM Base Map Attn Acc.$\uparrow$ Token Num.$\downarrow$ Reduction (%$\downarrow$
57.2 23,328 100
35.2 32 0.14%
R 38.9 134 0.57%
R 46.9 32 0.14%
R 37.7 32 0.14%
57.5 32 0.14%
Table 6: Ablation study on the impact of individual submodules on model performance. We incrementally add each proposed module component to the original baseline, following Sec. 3. A ✓indicates that the corresponding submodule is applied, while R denotes that its parameters are randomized. Note: For the ablation of w/. Token base + map, cross-dynamics attention parameters are randomized, as the vLLM cannot directly utilize the token dynamics map.
4.6 Ablation Study
Quantitative Ablation. To assess the contribution of each module, we conduct an ablation study by incrementally adding components and evaluating their impact. In Table 6, Base, Map, and Attn denote the token base (Sec. 3.3.2), token dynamics map (Sec. 3.3.3), and cross-dynamics attention (Sec. 3.4), respectively. Using only the Base reduces tokens to 0.14% of the original but results in a severe 22.0% accuracy drop, emphasizing the importance of spatial-temporal information. Adding the Map slightly improves accuracy but remains ineffective as the LLM cannot directly utilize motion cues. When cross-dynamics attention is introduced, spatial-temporal information is successfully integrated, achieving extreme token reduction while significantly restoring accuracy. To further verify module necessity, we randomize their parameters (R), leading to substantial accuracy drops, reinforcing their critical role in maintaining representation quality. This study highlights that while token clustering compresses redundant information, preserving motion dynamics via the token map and cross-dynamics attention is essential for maintaining accuracy in extreme token reduction settings.
Visualization of Token Clusters. As shown in Fig. 3, each object can be associated with one or multiple clusters and may be represented by either a single token or multiple tokens, as visual tokenizers act as descriptors based on similarity in token space. To analyze token clustering effectiveness, we compare an adaptive K-Means clustering approach with the state-of-the-art segmentation model Segment Anything (SAM) [20]. Specifically, adaptive K-Means is directly applied to token sequences, while the segmentation method assigns cluster indices based on the number of detected segmentation masks. While SAM achieves more refined clustering, both approaches successfully group similar objects and maintain cluster consistency across frames. Given the computational efficiency requirements of video large language models (vLLMs), adaptive K-Means provides a sufficiently precise clustering mechanism, effectively tracking object motions and trajectories for the subsequent token dynamics map.
Figure 3: Visualization of token clustering. Ideally, objects exhibiting similar visual features across frames should be consistently grouped into one or multiple clusters. We compare adaptive K-Means (with reduced cluster numbers for improved visualization) and Segment Anything as adaptive clustering methods.
5 Conclusion
In this work, we introduced Token Dynamics, a novel video representation framework designed to achieve extreme token reduction while preserving essential spatial-temporal coherence for video large language models (vLLMs). Our approach reformulates video representation by disentangling visual embeddings and motion information into two complementary components: (i) a concise token base that clusters object-descriptive tokens, and (ii) a token dynamics map that captures fine-grained spatial-temporal motion patterns. We further proposed a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, ensuring efficient and compact video representations. To systematically evaluate our framework, we introduced the novel Extreme Short Token Reduction task, encompassing both fixed-length and adaptive-length token compression settings. Our method demonstrated strong performance across multiple video-language benchmarks, reducing token count to as little as 0.07% of the original sequence while maintaining competitive accuracy. Specifically, we achieved only a minor 1.13% accuracy drop on the NextQA-MC benchmark and maintained comparable performance on ActNet-QA, Long Video Bench, and VideoMME, validating the effectiveness of our approach. Extensive ablation studies further highlighted the contributions of each module, confirming that both token base clustering and motion integration via cross-dynamics attention are critical for preserving video understanding. Moreover, efficiency analyses demonstrated the practical benefits of Token Dynamics, significantly reducing computational overhead compared to existing token reduction methods. Our framework is designed as a plug-and-play solution, making it broadly applicable to existing vLLMs without requiring architectural modifications. By balancing extreme token compression with high accuracy retention, Token Dynamics provides a scalable and efficient solution for video-language models, particularly in resource-constrained environments. Future work will explore further optimizations for large-scale models, extend token dynamics to more complex tasks like video generation and trajectory prediction.

### v2

@@
 To address this challenge, we propose Token Dynamics, a new video representation framework that dynamically reduces token count while preserving spatial-temporal coherence.
-Specifically, we disentangle video representations by separating visual embeddings from grid-level motion information, structuring them into: (i) a concise token base, created by clustering tokens that describe object-level content, and (ii) a token dynamics map, capturing detailed spatial-temporal motion patterns across grids.
-Furthermore, we introduce a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, thereby maintaining compactness and spatial-temporal integrity.
+Specifically, we disentangle video representations by separating visual embeddings from grid-level motion information, structuring them into: (i) a concise token hash table, created by clustering tokens that describe object-level content, (ii) a token indices key map, capturing detailed spatial-temporal motion patterns across grids, (iii) a token hash function, which vector-quantizes the token hash table to reconstruct the token sequence from the key map.
+Furthermore, we introduce a cross-dynamics attention mechanism that integrates motion features into the token hash table without increasing token length, thereby maintaining compactness and spatial-temporal integrity.
 We evaluate Token Dynamics on the proposed extreme short token reduction task, demonstrating a reduction of token count to merely 0.07% of the original tokens, with only a minor performance drop of 1.13% on the NextQA-MC benchmark.
@@
 Similarity-based merging techniques [3, 12, 13] often disregard spatial-temporal coherence by either directly merging positional encodings or reintroducing them post-merging, which impairs effective dynamic context representation.
-To address this, we introduce a structured token map representation, organized as Temporal $\times$ Width Coordinates $\times$ Height Dimension, to explicitly record spatial-temporal positions before clustering.
+To address this, we introduce a structured token indices key map representation, organized as Temporal $\times$ Width Coordinates $\times$ Height Dimension, to explicitly record spatial-temporal positions before clustering.
 To incorporate these dense maps without additional computational overhead, we propose a cross-dynamics attention mechanism that embeds motion information into clustered token sequences while maintaining their compact structure.
 Given video data’s inherent redundancy, tokens effectively represent grid-level visual information derived from pixels, offering a lightweight alternative to computationally intensive motion-tracking methods such as optical flow.
-Tracking token movements through the token map efficiently captures object grid-level dynamics, preserving essential spatial-temporal context with minimal computational cost.
+Tracking token movements through the token indices key map efficiently captures object grid-level dynamics, preserving essential spatial-temporal context with minimal computational cost.
 The third challenge lies in effectively evaluating concise token sequences.
@@
     We introduce a novel video representation framework featuring fixed/adaptive token clustering.
-It constructs a compact token base to capture visual pattern and object part descriptions and a token dynamics map to model their movements.
+It constructs a compact token hash table vector to capture visual pattern and object part descriptions and a token indices key map to model their movements.
 A cross-motion dynamics attention mechanism integrates spatial-temporal positional information without increasing token count.
@@
 Video Large Language Models (vLLMs) have emerged as powerful tools for bridging video understanding and natural language processing, enabling complex interpretations of video content through language-based interactions.
-Recent advancements have demonstrated remarkable capabilities in aligning visual and linguistic modalities, exemplified by frameworks such as LLaVA [15, 12, 14, 16], Flamingo [1], and MERLOT Reserve [29].
+Recent advancements have demonstrated remarkable capabilities in aligning visual and linguistic modalities, exemplified by frameworks such as LLaVA [16, 13, 15, 17], Flamingo [1], AuroraCap [5], and MERLOT Reserve [30].
 These methods typically rely on extensive pre-training using large-scale datasets like HD-VILA [27], InternVid [24], and NextQA [26], generating lengthy token sequences to represent videos effectively.
@@
 Such fixed strategies can either inadequately reduce redundant tokens, resulting in lingering inefficiencies, or inadvertently merge tokens representing distinct objects, thus losing crucial spatial-temporal dynamics necessary for precise video interpretation.
-To overcome these limitations, our proposed Token Dynamics framework introduces dynamic token clustering to generate a compact token representation while explicitly preserving spatial-temporal motion information via a dedicated token dynamics map.
+To overcome these limitations, our proposed Token Dynamics framework introduces dynamic token clustering to generate a compact token representation while explicitly preserving spatial-temporal motion information via a dedicated token indices key map.
 Through our novel cross-dynamics attention mechanism, our approach effectively integrates spatial-temporal coherence into concise token sequences without compromising accuracy, outperforming state-of-the-art token reduction methods, even in scenarios of extreme token compression.
@@
 The input video is first tokenized into a 1D token sequence and subsequently undergoes token clustering to group similar tokens, potentially representing the same object parts at the grid level.
-This clustering yields our proposed Token Dynamics representation, comprising two complementary components: (i) a concise token base, dynamically sized by merging tokens into representative cluster centroids, and (ii) a token dynamics map, a three-dimensional tensor explicitly capturing grid-level spatial-temporal coordinates (i.e., x, y, and frame index as spatial-temporal positions across frames) to represent detailed motion dynamics for each cluster.
-To efficiently integrate motion information into the token base, the token dynamics map is first flattened and dimensionally aligned via an MLP, and then processed by a cross-dynamics attention module employing two multi-head attention operations.
+This clustering yields our proposed Token Dynamics representation, comprising two complementary components: (i) a concise token hash table, dynamically sized by merging tokens into representative cluster centroids, and (ii) a token indices key map, a three-dimensional tensor explicitly capturing grid-level spatial-temporal coordinates (i.e., x, y, and frame index as spatial-temporal positions across frames) to represent detailed motion dynamics for each cluster.
+To efficiently integrate motion information into the token hash table, the token indices key map is first flattened and dimensionally aligned via an MLP, and then processed by a cross-dynamics attention module employing two multi-head attention operations.
 Crucially, this integration preserves the compactness of the original token count.
@@
 We conceptualize videos as compositions of multiple scenes, each containing objects with distinct attributes and motions.
-To achieve this, we reformulate the original video representation (token sequences from Vision Transformers [6]) into two components: (i) a concise token base, representing objects in each scene as token centroids obtained by clustering similar visual patterns at the grid-level; (ii) a token dynamics map, explicitly recording the spatial-temporal coordinates of each cluster centroid, capturing detailed object-level grid trajectories.
+To achieve this, we reformulate the original video representation (token sequences from Vision Transformers [7]) into two components: (i) a concise token hash table, representing objects in each scene as token centroids obtained by clustering similar visual patterns at the grid-level; (ii) a token indices key map, explicitly recording the spatial-temporal coordinates of each cluster centroid, capturing detailed object-level grid trajectories.
 As illustrated in Fig.
-1, this design ensures the token base and dynamics map are disentangled at the grid-level.
+1, this design ensures the token hash table and dynamics map are disentangled at the grid-level.
 Thus, we represent video content concisely without sacrificing crucial spatial-temporal information, achieving efficiency in token sequence reduction.
@@
 where $K$ is the number of clusters (fixed or adaptive), $\mathcal{F}_{\text{cluster}}$ denotes the clustering function (e.g., K-Means, Adaptive K-Means, or Video Segmentation [21]), $t_{i}$ represents the $i$-th token, $c_{i}$ is its assigned cluster index, and $s_{k}$ is the set of token indices assigned to the $k$-th cluster.
-3.3.2 Concise Token Base
-To construct a compact representation while preserving essential visual patterns and object part information, we merge clustered tokens into representative centroids, forming the concise token base.
-Using the token indices from Equation 2, the $k$-th token base $b_{k}$ is computed as:
+3.3.2 Concise Token Hash Table
+To construct a compact representation while preserving essential visual patterns and object part information, we merge clustered tokens into representative centroids, forming the concise token hash table.
+Using the token indices from Equation 2, the $k$-th token hash table $b_{k}$ vector is computed as:
 $b_{k}=\mathcal{M}(t,s_{k}),\quad k\in[1,K],\quad b\in\mathbb{R}^{K\times D}$
 where $b_{k}$ represents the centroid token for the $k$-th cluster, $s_{k}$ is the set of token indices assigned to the $k$-th cluster, and $\mathcal{M}$ denotes the merging function from Equation 1.
-3.3.3 Token Dynamics Map
-While the concise token base effectively represents visual patterns, it lacks spatial-temporal details essential for video understanding.
-To preserve motion dynamics, we record spatial-temporal coordinates during clustering, constructing a detailed motion representation—i.e., the token dynamics map.
+3.3.3 Token Indices Key Map
+While the concise token hash table effectively represents visual patterns, it lacks spatial-temporal details essential for video understanding.
+To preserve motion dynamics, we record spatial-temporal coordinates during clustering, constructing a detailed motion representation—i.e., the token indices key map.
 Specifically, we formulate a three-dimensional tensor capturing $x$, $y$, and frame indices, ensuring that each cluster centroid retains its original spatial-temporal position.
-This map aligns with the downsampled video structure from the ViT and is defined as:
+This map aligns with the downsampled video structure from the ViT and is defined by a token hash function:
 $m_{f,x,y}=c_{(fWH+xW+y)},\quad m\in\mathbb{R}^{T\times W\times H}$ (4)
@@
 3.4 Cross Dynamics Attention Module
-While the token base and token dynamics map effectively disentangle visual and motion information, directly embedding the token dynamics map would significantly increase computational overhead due to its large size.
-To address this, we propose a cross-dynamics attention module that integrates motion dynamics into the concise token base without increasing token count or dimensionality.
-First, the token dynamics map is projected via an MLP to align its dimensions with the token base.
-Then, a multi-head transformer decoder extracts motion information by treating the projected token dynamics map as encoded features and the token base as queries.
+While the token hash table and token indices key map effectively disentangle visual and motion information, directly embedding the token indices key map would significantly increase computational overhead due to its large size.
+To address this, we propose a cross-dynamics attention module that integrates motion dynamics into the concise token hash table without increasing token count or dimensionality.
+First, the token indices key map is projected via an MLP to align its dimensions with the token hash table.
+Then, a multi-head transformer decoder extracts motion information by treating the projected token indices key map as encoded features and the token hash table as queries.
 This process embeds spatial-temporal dynamics into the token sequence while maintaining its compact structure:
 $b_{\text{bank}}^{K\times D}=\mathcal{F}_{\text{A}}^{K\times D}\big{(}m^{T\times W\times H}W_{1}^{H\times D},\mathcal{F}_{\text{A}}^{K\times D}(b^{K\times D}W_{2}^{D\times D})\big{)}$
-where $\mathcal{F}_{\text{A}}$ represents multi-head attention, $m$ is the token dynamics map, $b$ is the concise token base, and $W_{r}$ denotes the weight matrix in the $r$-th MLP layer.
+where $\mathcal{F}_{\text{A}}$ represents multi-head attention, $m$ is the token indices key map, $b$ is the concise token hash table, and $W_{r}$ denotes the weight matrix in the $r$-th MLP layer.
 Superscripts indicate tensor shapes.
@@
 Note: For the ablation of w/.
-Token base + map, cross-dynamics attention parameters are randomized, as the vLLM cannot directly utilize the token dynamics map.
+Token Hash Table + map, cross-dynamics attention parameters are randomized, as the vLLM cannot directly utilize the token indices key map.
 4.6 Ablation Study
@@
 To assess the contribution of each module, we conduct an ablation study by incrementally adding components and evaluating their impact.
-In Table 6, Base, Map, and Attn denote the token base (Sec.
-3.3.2), token dynamics map (Sec.
+In Table 6, Base, Map, and Attn denote the token hash table (Sec.
+3.3.2), token indices key map (Sec.
 3.3.3), and cross-dynamics attention (Sec.
@@
 While SAM achieves more refined clustering, both approaches successfully group similar objects and maintain cluster consistency across frames.
-Given the computational efficiency requirements of video large language models (vLLMs), adaptive K-Means provides a sufficiently precise clustering mechanism, effectively tracking object motions and trajectories for the subsequent token dynamics map.
+Given the computational efficiency requirements of video large language models (vLLMs), adaptive K-Means provides a sufficiently precise clustering mechanism, effectively tracking object motions and trajectories for the subsequent token indices key map.
 Figure 3: Visualization of token clustering.
@@
 In this work, we introduced Token Dynamics, a novel video representation framework designed to achieve extreme token reduction while preserving essential spatial-temporal coherence for video large language models (vLLMs).
-Our approach reformulates video representation by disentangling visual embeddings and motion information into two complementary components: (i) a concise token base that clusters object-descriptive tokens, and (ii) a token dynamics map that captures fine-grained spatial-temporal motion patterns.
-We further proposed a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, ensuring efficient and compact video representations.
+Our approach reformulates video representation by disentangling visual embeddings and motion information into two complementary components: (i) a concise token hash table that clusters object-descriptive tokens, and (ii) a token indices key map that captures fine-grained spatial-temporal motion patterns.
+We further proposed a cross-dynamics attention mechanism that integrates motion features into the token hash table without increasing token length, ensuring efficient and compact video representations.
 To systematically evaluate our framework, we introduced the novel Extreme Short Token Reduction task, encompassing both fixed-length and adaptive-length token compression settings.
@@
 Specifically, we achieved only a minor 1.13% accuracy drop on the NextQA-MC benchmark and maintained comparable performance on ActNet-QA, Long Video Bench, and VideoMME, validating the effectiveness of our approach.
-Extensive ablation studies further highlighted the contributions of each module, confirming that both token base clustering and motion integration via cross-dynamics attention are critical for preserving video understanding.
+Extensive ablation studies further highlighted the contributions of each module, confirming that both token hash table clustering and motion integration via cross-dynamics attention are critical for preserving video understanding.
 Moreover, efficiency analyses demonstrated the practical benefits of Token Dynamics, significantly reducing computational overhead compared to existing token reduction methods.